If a 60 m^3 of water are pumped per hour to a height of 80m, calculate a) the power required, in kilowatts, and (b) the work done, in mega joules, in 10 minutes. Neglect any losses and assume that 1 m^3 of water has a mass of 1000kg.
1
Expert's answer
2018-12-07T10:11:10-0500

The power P= E/t, where t = 1 hour = 3600 s, E=mgh , then P=(1000 kg ∙60〖 m〗^3∙10 m/s^2 ∙80 m)/(3600 s)=13.3×〖10〗^3 W=13.3 kW

the work done in 10 minutes A=Pt=13.3×〖10〗^3 W∙600 s=8×〖10〗^6 J=8 MJ